evening shift 8/22/99

John Shields (jshields@ksera)
Sun, 22 Aug 1999 23:57:18 -0500

Shift Summary for eve shift, 08/22/99

On Shift: J. Shields, J. Ladue

Hours of beam available at all intensities: 8.3
Hours of beam available at nominal intensity: 8.1
Hours of E832 data (special runs not included): 8.1
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

booster vacuum burst, tape changes

========================================================================
Runs ended this shift:
Run 14300 - 23.7M events - BEAM_832

Tapes brought to FCC this shift:
(including tapes for online split)

on-line spilt: QKNL25
========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:
E832 datataking, brief (~10 min) beam shut down, fixed a red SDAQ
button, and swapped some tapes.

In short: If ya weren't here, ya didn't miss much.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14300 155 15:58 ktevana errors: missing FFFF word stream 5, RWP9.
Kfrended. Problem solved.

185 16:42 Beam goes away due to a Booster Vacuum Burst.
Estimate is 10 min.

195 16:51 Amazingly enough, beam returns as promised.

233 17:43 ktevana errors: unknown crate, stream 5. Once
again, kfrend saves the day.

18:00 Mounted on-line split tape QKNL25

304 19:18 ktevana errors: missing FFFF word stream 5, RWP9.
Kfrended.

330 20:05 red button: SDAQ. JL re-runs ktsad.exe
Problem solved.

20:25 Run stops. Tapes full. Mounted tapes
QK*312-14 and QKB109

20:34 Mount QKLA14 for on-split

14301 20:41 Began new run

20:54 SEM consistantly > 7.0E12. Shift crew calls MCR
& asks them to reduce it a little. The SEM
soon drifts back down to nominal level.

20:55 DC Histograms screen goes blank. Ran fix-disp,
so next shift crew needs to drt_fresh_start
when they initialize the next run.

21:24 "Val Alarm" goes off. CSI chan 1137 has laser
stability of 10.4 for 2 spills. It then
returns to normal.

23:17 ktevana errors: missing FFFF word streams, RWP9.
Kfrended.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

do a drt_fresh_start before you initialize the next run